# A candidate proton cyclotron feature in the ultraluminous X-ray source NGC 4656 ULX-1

### 11 Mar 2026 v1

@@
Ultraluminous X-ray sources represent extreme super-Eddington accretion regimes, and a subset is now known to host highly magnetized neutron stars. However, direct observational probes of their surface magnetic fields remain scarce. In this Letter, we report the detection of a narrow X-ray absorption feature at keV in the XMM–Newton/EPIC-pn spectrum of NGC 4656 ULX-1. The source exhibits a hard-ultraluminous state, while our timing analysis reveals a candidate pulsation at 0.9736 Hz, with a local significance of and a pulsed fraction of . The feature is robust against changes in continuum modeling and data-selection criteria, retaining a statistical significance of in Monte Carlo simulations. Interpreting the absorption as a proton cyclotron resonant scattering feature implies a local magnetic field of – G in the line-forming region. This value is consistent with strong magnetic fields anchored near the neutron star surface, even if the large-scale dipole is substantially weaker. Although we discuss electron cyclotron features and atomic transitions as possible alternatives, they appear less consistent with the observed phenomenology.
@@
Ultraluminous X-ray sources (ULXs) are off-nuclear objects with apparent luminosities of – erg s-1, exceeding the Eddington limit for typical stellar-mass compact objects (Feng and Soria 2011; Kaaret et al. 2017; King et al. 2023). The consensus attributes this extreme output to super-Eddington accretion onto stellar-mass black holes or neutron stars, facilitated by geometric beaming (Poutanen et al. 2007; King 2009) and strong radiation-driven winds (Pinto et al. 2016; Kosec et al. 2018, 2021). The discovery of coherent X-ray pulsations in several ULXs definitively established that a subset of the population harbors neutron stars (Bachetti et al. 2014; Fürst et al. 2016; Israel et al. 2017; Carpano et al. 2018; Walton et al. 2018).
Cyclotron resonant scattering features (CRSFs) provide one of the few direct probes of these neutron-star magnetic fields (Mészáros 1992; Caballero and Wilms 2012; Staubert et al. 2019). In Galactic X-ray pulsars, electron CRSFs at tens of keV commonly indicate surface magnetic fields of order G. These features often display harmonics and a strong pulse-phase dependence (Trümper et al. 1978; Coburn et al. 2002; Staubert et al. 2019). In contrast, proton cyclotron features occur at energies lower by a factor of 1836 than electron cyclotron lines, reflecting the proton-to-electron mass ratio, and for magnetar-strength fields are therefore expected at a few keV. A narrow proton absorption line at a few keV would imply a magnetar-strength field ( G) at the neutron star surface (Ho and Lai 2003; Harding and Lai 2006). These considerations motivate phase- and state-resolved searches for CRSFs in ULXs, where Comptonized accretion columns and scattering in optically thick winds can significantly shape the observed spectra (Becker and Wolff 2007; Middleton et al. 2015; Walton et al. 2018).
@@ 3.2 Spectral analysis
The 0.3–10 keV EPIC-pn spectrum is well described by a simple absorbed powerlaw model (), though the residuals reveal a clear, narrow absorption feature near keV. Adding a multiplicative Gaussian absorption component (gabs) improves the fit by for three additional free parameters. Because standard likelihood-ratio tests (LRT) can overestimate the significance of spectral-line components, all detection significances were calibrated via Monte Carlo LRT simulations (Protassov et al. 2002). Comparing the null powerlaw continuum against the line-included model (Fig. 5) yields () for pn alone. Including the MOS2 spectrum lowers this feature detectability to (); the MOS2 data alone show no significant feature (), consistent with their lower effective area near 3.3 keV.
@@
For standard ULX spectral classification, the canonical diskbb+powerlaw continuum (Sutton et al. 2013) was applied to the joint XMM–Newton (pn+MOS2) + NuSTAR (FPMA/B) broadband dataset over 0.3–25 keV (allowing for cross-calibration constants), yielding (), with keV and , placing NGC 4656 ULX-1 firmly in the hard-ultraluminous (HUL) regime. Adopting a distance of 8.21 Mpc (Muñoz-Elgueta et al. 2018) for consistency with recent optical and spectroscopic studies of the host galaxy, this corresponds to an unabsorbed 0.3–25 keV luminosity of . The Gaussian absorption component at 3.3 keV improves the fit by , corresponding to a detection significance of using Monte Carlo simulations.
A physically motivated thermally Comptonized model, ThComp(diskbb), gives a comparable improvement (, ), with the electron temperature poorly constrained ( keV) within the available energy range. The line parameters remain stable across all continua tested, confirming that the feature is not an artefact of the continuum choice.
The best-fit parameters for both two-component continuum models (with and without the gabs line) are summarised in Table 1. Figure 1 shows the unfolded broadband spectra and residuals for the diskbb+powerlaw case. To assess whether non-detections in MOS2 and FPMA/B artificially bias the line significance, we re-evaluated the joint fit ignoring the 3.1–3.4 keV bins for these instruments; the EPIC-pn data in this band then yield a revised significance of .
@@ 4 Discussion and Conclusions
The timing analysis provides additional context, though it is still tentative. A coherent pulsation search in the XMM–Newton EPIC-pn data identifies a possible candidate signal at Hz that reaches a local significance of and corresponds to a pulsed fraction of . The signal is most prominent below 3 keV and is not detected in the simultaneous NuSTAR observation. Phase-resolved spectra extracted by folding the event list on this candidate period suggest that the absorption feature may be somewhat stronger during the pulse-off phase. However, the current statistics are limited, and deeper observations are needed to confirm this dependence. Given the tentative nature of the timing evidence, we first assess whether the feature could have an atomic origin (see Appendix A.3).
ULXs frequently show highly ionized, fast outflows that produce narrow blueshifted absorption features (Pinto et al. 2016; Kosec et al. 2018, 2021). Among atomic transitions in this energy range, the closest match to our measured centroid is S xvi at a rest energy of keV (van Hoof 2018). Interpreting the feature in this way would require a blueshift of km s-1 to reproduce the observed centroid at keV (see Table 1), a value broadly consistent with wind velocities expected in this class of systems.
@@
These observations suggest that a proton CRSF offers a more plausible physical interpretation (Ho and Lai 2003; Harding and Lai 2006). For a fundamental proton cyclotron line, , the observed centroid implies a local magnetic field – G in the line-forming region for gravitational redshifts –, i.e. in the magnetar regime. Strictly speaking, this estimate refers to the near-surface field strength where the feature is produced and does not necessarily coincide with the large-scale magnetic dipole strength.
@@
In the HUL geometry, our line of sight likely peers into the accretion funnel, giving an unobstructed view of the accretion column and reduced wind covering (Middleton et al. 2015). Under this scenario, the line centroid appears likely to remain stable (set by the magnetic field strength ), while the equivalent width may vary with accretion rate or geometry (e.g. between HUL and soft-ultraluminous states). Higher harmonics would be intrinsically weak or absent in the proton case, as observed. The candidate pulsed fraction of is also consistent with this picture.
@@ Acknowledgements.
- XSPEC: the first ten years. In Astronomical Data Analysis Software and Systems V, G. H. Jacoby and J. Barnes (Eds.), ASP Conference Series, Vol. 101, San Francisco, pp. 17. Cited by: §3.2.
- An ultraluminous x-ray source powered by an accreting neutron star. Nature 514 (7521), pp. 202–204. External Links: Document Cited by: §B.2, §1.
- HENDRICS: High ENergy Data Reduction Interface from the Command Shell. Note: Astrophysics Source Code Library, record ascl:1805.019 External Links: 1805.019, ADS entry Cited by: §A.2, §B.2, §3.1.
- Thermal and bulk comptonization in accretion-powered x-ray pulsars. The Astrophysical Journal 654 (1), pp. 435–457. External Links: Document Cited by: §1.
@@
- Cyclotron lines in x-ray pulsars: a brief review. Memorie della Società Astronomica Italiana 83, pp. 230–237. Cited by: §1.
- Discovery of pulsations from ngc 300 ulx1 and its fast period evolution. Monthly Notices of the Royal Astronomical Society: Letters 476 (1), pp. L45–L49. External Links: Document Cited by: §1.
- Magnetic fields of accreting x-ray pulsars with the rossi x-ray timing explorer. The Astrophysical Journal 580 (1), pp. 394–412. External Links: Document Cited by: §1.
- A broadband X-ray view of the ultraluminous sources in the galaxies ESO 501–023 and IC 5052. A&A 693, pp. A122. External Links: Document, ADS entry Cited by: §A.2.
- The hard ultraluminous state of NGC 5055 ULX X-1. arXiv e-prints, pp. arXiv:2511.13686. External Links: Document, 2511.13686, ADS entry Cited by: §A.2.
- The h-test probability distribution revisited: improved sensitivity. Astronomy and Astrophysics 517, pp. L9. External Links: ISSN 1432-0746, Link, Document Cited by: §A.2.
- Ultraluminous x-ray sources in the chandra and xmm-newton era. New Astronomy Reviews 55 (5), pp. 166–183. External Links: Document Cited by: §1.
@@
- Discovery of coherent pulsations from the ultraluminous x-ray source ngc 7793 p13. The Astrophysical Journal Letters 831 (2), pp. L14. External Links: Document Cited by: §1.
- Physics of strongly magnetized neutron stars. Reports on Progress in Physics 69 (9), pp. 2631–2708. External Links: Document Cited by: §1, §4, §4.
- Atmospheres and spectra of strongly magnetized neutron stars: ii. the effect of vacuum polarization. Monthly Notices of the Royal Astronomical Society 338 (1), pp. 233–252. External Links: Document Cited by: §1, §4, §4.
@@
- Discovery of cyclotron resonance features in the burst spectra of sgr 1806-20. The Astrophysical Journal Letters 574, pp. L51–L55. External Links: Document Cited by: §1.
@@
- Ultraluminous x-ray sources. Annual Review of Astronomy and Astrophysics 55, pp. 303–341. External Links: Document Cited by: §1.
- ULXs: neutron stars versus black holes. Monthly Notices of the Royal Astronomical Society: Letters 393 (1), pp. L41–L44. External Links: Document Cited by: §1.
- Ultraluminous x-ray sources. New Astronomy Reviews 96, pp. 101672. External Links: ISSN 1387-6473, Link, Document Cited by: §1.
- Ultrafast outflows in ultraluminous x-ray sources. Monthly Notices of the Royal Astronomical Society 479 (3), pp. 3978–3994. External Links: Document Cited by: §1, §4, §4.
- High-resolution x-ray spectroscopy of ultraluminous x-ray sources: winds and accretion geometry. Monthly Notices of the Royal Astronomical Society 501 (2), pp. 2412–2432. External Links: Document Cited by: §1, §4, §4.
- Atmospheres and spectra of strongly magnetized neutron stars ii: effect of vacuum polarization. The Astrophysical Journal 566 (1), pp. 373–377. External Links: Document Cited by: §4.
@@
- Diagnosing the accretion flow in ultraluminous x-ray sources with x-ray spectral and variability analysis. Monthly Notices of the Royal Astronomical Society 454 (3), pp. 3134–3148. External Links: Document Cited by: §1, §4, §4.
- Kinematics and physical properties of the nearby galaxy ngc 4656 and its tdg candidate. Monthly Notices of the Royal Astronomical Society 480 (3), pp. 3257–3278. External Links: ISSN 0035-8711, Document, Link, https://academic.oup.com/mnras/article-pdf/480/3/3257/25519360/sty1934.pdf Cited by: §1, §3.2.
@@
- Supercritically accreting stellar mass black holes as ultraluminous x-ray sources. Monthly Notices of the Royal Astronomical Society 377 (3), pp. 1187–1194. External Links: Document Cited by: §1.
- Statistics, handle with care: detecting multiple model components with the likelihood ratio test. The Astrophysical Journal 571, pp. 545–559. External Links: Document Cited by: §3.2.
- A variable absorption feature at 8.1 kev in the x-ray spectrum of the anomalous x-ray pulsar 1rxs j170849.0-400910. The Astrophysical Journal Letters 586 (2), pp. L65–L68. External Links: Document Cited by: §1.
@@
- Discovery of absorption features in the x-ray spectrum of an isolated neutron star. The Astrophysical Journal Letters 574 (1), pp. L61–L64. External Links: Document Cited by: §1.
- From spitzer galaxy photometry to tully–fisher distances. Monthly Notices of the Royal Astronomical Society 444 (1), pp. 527–541. External Links: ISSN 0035-8711, Document, Link, https://academic.oup.com/mnras/article-pdf/444/1/527/18753576/stu1450.pdf Cited by: §1.
- Cyclotron lines in highly magnetized neutron stars. Astronomy & Astrophysics 622, pp. A61. External Links: Document Cited by: §1, §4.
@@
- Evidence for strong cyclotron line emission in the hard x-ray spectrum of hercules x-1. The Astrophysical Journal Letters 219, pp. L105–L110. External Links: Document Cited by: §1.
@@
- Super-eddington accretion on to the neutron star ngc 7793 p13. Monthly Notices of the Royal Astronomical Society 473 (4), pp. 4360–4370. External Links: Document Cited by: §1, §1.
- On the absorption of x-rays in the interstellar medium. The Astrophysical Journal 542 (2), pp. 914–924. External Links: Document Cited by: §3.2.

### 12 Mar 2026 v2

@@
Ultraluminous X-ray sources represent extreme super-Eddington accretion regimes, and a subset is now known to host highly magnetized neutron stars. However, direct observational probes of their surface magnetic fields remain scarce. In this Letter, we report the detection of a narrow X-ray absorption feature at keV in the XMM–Newton/EPIC-pn spectrum of NGC 4656 ULX-1. The source exhibits a hard ultraluminous state, while our timing analysis reveals a candidate pulsation at 0.9736 Hz, with a local significance of and a pulsed fraction of . The feature is robust against changes in continuum modeling and data-selection criteria, retaining a statistical significance of in Monte Carlo simulations. Interpreting the absorption as a proton cyclotron resonant scattering feature implies a local magnetic field of – G in the line-forming region. This value is consistent with strong magnetic fields anchored near the neutron star surface, even if the large-scale dipole is substantially weaker. Although we discuss electron cyclotron features and atomic transitions as possible alternatives, they appear to be less consistent with the observed phenomenology.
@@
Ultraluminous X-ray sources (ULXs) are off-nuclear objects with apparent luminosities of – erg s-1, exceeding the Eddington limit for typical stellar-mass compact objects (Feng and Soria 2011; Kaaret et al. 2017; King et al. 2023). The consensus attributes this extreme output to super-Eddington accretion onto stellar-mass black holes or neutron stars, facilitated by geometric beaming (Poutanen et al. 2007; King 2009) and strong radiation-driven winds (Pinto et al. 2016; Kosec et al. 2018, 2021). The discovery of coherent X-ray pulsations in several ULXs definitively established that a subset of the population harbors neutron stars (Bachetti2014; Fürst et al. 2016; Israel et al. 2017; Carpano et al. 2018; Walton et al. 2018).
Cyclotron resonant scattering features (CRSFs) are among the few direct probes of these neutron-star magnetic fields (Mészáros 1992; Caballero and Wilms 2012; Staubert et al. 2019). In Galactic X-ray pulsars, electron CRSFs at tens of kilo-electronvolts commonly indicate surface magnetic fields of order G. These features often display harmonics and a strong pulse-phase dependence (Trümper et al. 1978; Coburn et al. 2002; Staubert et al. 2019). In contrast, proton cyclotron features occur at energies lower by a factor of 1836 than electron cyclotron lines, reflecting the proton-to-electron mass ratio, and for magnetar-strength fields are therefore expected at a few kilo-electronvolts. A narrow proton absorption line at a few kilo-electronvolts would imply a magnetar-strength field ( G) at the neutron star surface (Ho and Lai 2003; Harding and Lai 2006). These considerations motivate phase- and state-resolved searches for CRSFs in ULXs, where Comptonized accretion columns and scattering in optically thick winds can significantly shape the observed spectra (Becker and Wolff 2007; Middleton et al. 2015; Walton et al. 2018).
@@ 3.2 Spectral analysis
The 0.3–10 keV EPIC-pn spectrum is well described by a simple absorbed powerlaw model (), though the residuals reveal a clear, narrow absorption feature near keV. Adding a multiplicative Gaussian absorption component (gabs) improves the fit by for three additional free parameters. Because standard likelihood-ratio tests (LRTs) can overestimate the significance of spectral-line components, all detection significances were calibrated via Monte Carlo LRT simulations (Protassov et al. 2002). Comparing the null powerlaw continuum against the line-included model (Fig. 5) yields () for pn alone. Including the MOS2 spectrum lowers this feature detectability to (); the MOS2 data alone show no significant feature (), consistent with their lower effective area near 3.3 keV.
@@
For standard ULX spectral classification, the canonical diskbb+powerlaw continuum (Sutton et al. 2013) was applied to the joint XMM–Newton (pn+MOS2) + NuSTAR (FPMA/B) broadband dataset over 0.3–25 keV (allowing for cross-calibration constants), yielding (), with keV and , placing NGC 4656 ULX-1 firmly in the hard ultraluminous (HUL) regime. Adopting a distance of 8.21 Mpc (Muñoz-Elgueta et al. 2018) for consistency with recent optical and spectroscopic studies of the host galaxy, this corresponds to an unabsorbed 0.3–25 keV luminosity of . The Gaussian absorption component at 3.3 keV improves the fit by , corresponding to a detection significance of using Monte Carlo simulations.
A physically motivated thermally Comptonized model, ThComp(diskbb), gives a comparable improvement (, ), with the electron temperature poorly constrained ( keV) within the available energy range. The line parameters remain stable across all continua tested, confirming that the feature is not an artifact of the continuum choice.
The best-fit parameters for both two-component continuum models (with and without the gabs line) are summarized in Table 1. Figure 1 shows the unfolded broadband spectra and residuals for the diskbb+powerlaw case. To assess whether non-detections in MOS2 and FPMA/B artificially bias the line significance, we reevaluated the joint fit ignoring the 3.1–3.4 keV bins for these instruments; the EPIC-pn data in this band then yield a revised significance of .
@@ 4 Discussion and conclusions
The timing analysis provides additional context, though it is still tentative. A coherent pulsation search in the XMM–Newton EPIC-pn data identifies a possible candidate signal at Hz that reaches a local significance of and corresponds to a pulsed fraction of . The signal is most prominent below 3 keV and is not detected in the simultaneous NuSTAR observation. Phase-resolved spectra extracted by folding the event list on this candidate period suggest that the absorption feature may be somewhat stronger during the pulse-off phase. However, the current statistics are limited, and deeper observations are needed to confirm this dependence. Given the tentative nature of the timing evidence, we first assessed whether the feature could have an atomic origin (see Appendix A.3).
Ultraluminous X-ray sources frequently show highly ionized, fast outflows that produce narrow blueshifted absorption features (Pinto et al. 2016; Kosec et al. 2018, 2021). Among atomic transitions in this energy range, the closest match to our measured centroid is S xvi at a rest energy of keV (van Hoof 2018). Interpreting the feature in this way would require a blueshift of km s-1 to reproduce the observed centroid at keV (see Table 1), a value broadly consistent with wind velocities expected in this class of systems.
@@
These observations suggest that a proton CRSF offers a more plausible physical interpretation (Ho and Lai 2003; Harding and Lai 2006). For a fundamental proton cyclotron line, , the observed centroid implies a local magnetic field of – G in the line-forming region for gravitational redshifts –, i.e., in the magnetar regime. Strictly speaking, this estimate refers to the near-surface field strength where the feature is produced and does not necessarily coincide with the large-scale magnetic dipole strength.
@@
In the HUL geometry, our line of sight likely peers into the accretion funnel, giving an unobstructed view of the accretion column and reduced wind covering (Middleton et al. 2015). Under this scenario, the line centroid appears likely to remain stable (set by the magnetic field strength ), while the equivalent width may vary with the accretion rate or geometry (e.g. between HUL and soft ultraluminous states). Higher harmonics would be intrinsically weak or absent in the proton case, as is observed. The candidate pulsed fraction of is also consistent with this picture.
@@ Acknowledgements.
- XSPEC: The First Ten Years. In Astronomical Data Analysis Software and Systems V, G. H. Jacoby and J. Barnes (Eds.), Astronomical Society of the Pacific Conference Series, Vol. 101, pp. 17. Cited by: §3.2.
- An ultraluminous x-ray source powered by an accreting neutron star. Nature 514 (7521), pp. 202–204. External Links: ISSN 1476-4687, Link, Document Cited by: §B.2.
- HENDRICS: High ENergy Data Reduction Interface from the Command Shell. Note: Astrophysics Source Code Library, ascl:1805.019 External Links: 1805.019 Cited by: §A.2, §B.2, §3.1.
- Thermal and Bulk Comptonization in Accretion-powered X-Ray Pulsars. ApJ 654 (1), pp. 435–457. External Links: Document, astro-ph/0609035 Cited by: §1.
@@
- Cyclotron lines in x-ray pulsars: a brief review. Mem. Soc. Astron. Italiana 83, pp. 230–237. Cited by: §1.
- Discovery of pulsations from ngc 300 ulx1 and its fast period evolution. MNRAS: Letters 476 (1), pp. L45–L49. External Links: Document Cited by: §1.
- Magnetic fields of accreting x-ray pulsars with the rossi x-ray timing explorer. ApJ 580 (1), pp. 394–412. External Links: Document Cited by: §1.
- A broadband x-ray view of the ultraluminous sources in the galaxies eso 501–023 and ic 5052. A&A 693, pp. A122. External Links: Document, Link Cited by: §A.2.
- The hard ultraluminous state of ngc 5055 ulx x-1. A&A 705, pp. A136. External Links: Document, Link Cited by: §A.2.
- The h-test probability distribution revisited: improved sensitivity. A&A 517, pp. L9. External Links: ISSN 1432-0746, Link, Document Cited by: §A.2.
- Ultraluminous x-ray sources in the chandra and xmm-newton era. New Astron. Rev. 55 (5), pp. 166–183. External Links: Document Cited by: §1.
@@
- Discovery of coherent pulsations from the ultraluminous x-ray source ngc 7793 p13. ApJ 831 (2), pp. L14. External Links: Document Cited by: §1.
- Physics of strongly magnetized neutron stars. Rep. Prog. Phys. 69 (9), pp. 2631–2708. External Links: Document Cited by: §1, §4, §4.
- Atmospheres and spectra of strongly magnetized neutron stars: ii. the effect of vacuum polarization. MNRAS 338 (1), pp. 233–252. External Links: Document Cited by: §1, §4, §4.
@@
- Discovery of cyclotron resonance features in the burst spectra of sgr 1806-20. ApJL 574, pp. L51–L55. External Links: Document Cited by: §1.
@@
- Ultraluminous x-ray sources. ARA&A 55, pp. 303–341. External Links: Document Cited by: §1.
- ULXs: neutron stars versus black holes. MNRAS 393 (1), pp. L41–L44. External Links: Document Cited by: §1.
- Ultraluminous x-ray sources. New Astron. Rev. 96, pp. 101672. External Links: ISSN 1387-6473, Link, Document Cited by: §1.
- Ultrafast outflows in ultraluminous x-ray sources. MNRAS 479 (3), pp. 3978–3994. External Links: Document Cited by: §1, §4, §4.
- High-resolution x-ray spectroscopy of ultraluminous x-ray sources: winds and accretion geometry. MNRAS 501 (2), pp. 2412–2432. External Links: Document Cited by: §1, §4, §4.
- Atmospheres and spectra of strongly magnetized neutron stars ii: effect of vacuum polarization. ApJ 566 (1), pp. 373–377. External Links: Document Cited by: §4.
@@
- Diagnosing the accretion flow in ultraluminous x-ray sources with x-ray spectral and variability analysis. MNRAS 454 (3), pp. 3134–3148. External Links: Document Cited by: §1, §4, §4.
- Kinematics and physical properties of the nearby galaxy ngc 4656 and its tdg candidate. MNRAS 480 (3), pp. 3257–3278. External Links: ISSN 0035-8711, Document, Link, https://academic.oup.com/mnras/article-pdf/480/3/3257/25519360/sty1934.pdf Cited by: §1, §3.2.
@@
- Supercritically accreting stellar mass black holes as ultraluminous x-ray sources. MNRAS 377 (3), pp. 1187–1194. External Links: Document Cited by: §1.
- Statistics, handle with care: detecting multiple model components with the likelihood ratio test. ApJ 571, pp. 545–559. External Links: Document Cited by: §3.2.
- A variable absorption feature at 8.1 kev in the x-ray spectrum of the anomalous x-ray pulsar 1rxs j170849.0-400910. ApJL 586 (2), pp. L65–L68. External Links: Document Cited by: §1.
@@
- Discovery of absorption features in the x-ray spectrum of an isolated neutron star. ApJL 574 (1), pp. L61–L64. External Links: Document Cited by: §1.
- From spitzer galaxy photometry to tully–fisher distances. MNRAS 444 (1), pp. 527–541. External Links: ISSN 0035-8711, Document, Link, https://academic.oup.com/mnras/article-pdf/444/1/527/18753576/stu1450.pdf Cited by: §1.
- Cyclotron lines in highly magnetized neutron stars. A&A 622, pp. A61. External Links: Document Cited by: §1, §4.
@@
- Evidence for strong cyclotron line emission in the hard x-ray spectrum of hercules x-1. ApJL 219, pp. L105–L110. External Links: Document Cited by: §1.
@@
- Super-eddington accretion on to the neutron star ngc 7793 p13. MNRAS 473 (4), pp. 4360–4370. External Links: Document Cited by: §1, §1.
- On the absorption of x-rays in the interstellar medium. ApJ 542 (2), pp. 914–924. External Links: Document Cited by: §3.2.